If Cosa=kcosb then show that tan(a+b/2)=(1-k)/(1+k)cot(a-b/2)

Grade:11

1 Answers

sonika p
6 years ago
cosa/cosb = k
(cos b – cos a)/(cos b + cos a)  = (1-k) / (1+k)
[sin ((a-b)/2)sin((a+b)/2)]/ [cos ((a-b)/2)cos((a+b)/2)] = (1-k) / (1+k)
[tan ((a-b)/2)tan((a+b)/2)] = (1-k) / (1+k)
tan((a+b)/2) = [(1-k) / (1+k)] cot ((a-b)/2)